# Through the Looking Glass: Mirror Schrödinger Bridges

### v1

@@ 4.1 Alternating Minimization Procedure
Take the reference path measure ${\mathbb{P}}^{0}$ to be time-symmetric. As an example, we can take ${\mathbb{P}}^{0}$ to be associated to an Ornstein–Uhlenbeck process ${\mathbf{X}}_{t}$ given by an SDE of the form ${\rm{d}}{\mathbf{X}}_{t}=-\alpha{\mathbf{X}}_{t}{\rm{d}}t+\sigma{\rm{d}}{\mathbf{W}}_{t}$, for $\alpha>0$, or more generally any reversible diffusion process. We propose the following iterative scheme:
${\mathbb{P}}^{2k+1}=\operatorname*{arg\,min}_{{\mathbb{P}}\in{\mathbb{D}}(\pi,\cdot)}D_{\mathrm{KL}}\left({\mathbb{P}}\;\|\;{\mathbb{P}}^{2k}\right)$ (direct $D_{\mathrm{KL}}$ projection) (4)
${\mathbb{P}}^{2k+2}=\operatorname*{arg\,min}_{{\mathbb{P}}\in{\mathbb{S}}}D_{\mathrm{KL}}\big{(}{{\mathbb{P}}^{2k+1}}\;\|\;{\mathbb{P}}\big{)},$ (reverse $D_{\mathrm{KL}}$ projection) (5)
@@
where ${\mathbb{S}}$ is the set of time-symmetric path measures with no marginal constraints. This scheme is an instance of AMP and differs from IPFP in that it alternates between direct and reverse $D_{\mathrm{KL}}$ projections. To see this, note that (4) is a direct $D_{\mathrm{KL}}$ projection and coincides with the odd-numbered steps in the IPFP iterations (2), whereas (5) is a reverse $D_{\mathrm{KL}}$ projection, as the KL divergence is being computed against the optimization parameter ${\mathbb{P}}$ instead of the previously produced path measure ${\mathbb{P}}^{2k+1}$. That is to say, each iteration of the AMP scheme in (4)-(5) is designed to obtain the time-symmetric measure ${\mathbb{P}}$ that minimizes the objective while remaining close in KL divergence $D_{\mathrm{KL}}$ to the measure obtained in the previous half iteration, which satisfies the initial marginal constraint $\pi$.
It is natural to ask why we consider reverse $D_{\mathrm{KL}}$ projections, as opposed to direct projections, onto the space of symmetric path measures. In fact, replacing (5) with a direct $D_{\mathrm{KL}}$ projection would result in a viable symmetrized variant of IPFP, and by (Ruschendorf, 1995), the resulting iterates would converge in total variation to the mirror Schrödinger bridge. The difficulty is in computing the direct $D_{\mathrm{KL}}$ projection of a path measure onto ${\mathbb{S}}$. As we will demonstrate in section 4.3, it is considerably easier to compute the reverse $D_{\mathrm{KL}}$ projection onto ${\mathbb{S}}$, as this particular projection can be done completely analytically.
4.2 Convergence
For the scheme in steps (4)-(5) to be practical, we must prove that the iterates ${\mathbb{P}}^{k}$ converge to the mirror Schrödinger bridge ${\mathbb{P}}_{\rm{MSB}}$. The pointwise convergence of schemes like steps (4)-(5) was established by Csiszár & Tusnády (1984) in the special case where the state space is finite. In our setting, however, we work with infinite state spaces of the form $\mathbb{R}^{n}$ for some dimension $n>0$. In the following theorem, we prove that the sequence obtained in the AMP scheme converges in total variation to the mirror Schrödinger bridge, without relying on the finiteness assumption for the state space. To our knowledge, this result has not been established previously in the literature.
Theorem 1.
Let ${\mathbb{P}}^{k}$ be the sequence of path measures obtained via the alternating minimization procedure defined in steps ((((4))))-((((5)))). Then ${\mathbb{P}}^{k}$ converges to ${\mathbb{P}}_{\rm{MSB}}$ in total variation as $k\to\infty$.
Our proof strategy for Theorem 1 is inspired by the convergence proofs for IPFP given in (Ruschendorf, 1995, Proposition 2.1) and (De Bortoli et al., 2021, Theorem 36). The basic idea is to prove that the sequence ${\mathbb{P}}^{k}$ is Cauchy with respect to the metric $\delta_{\rm{TV}}$ induced by total variation; we then conclude using completeness of the space of path measures together with optimality of the Schrödinger bridge. The crucial distinction between our setting and theirs is that one of our $D_{\mathrm{KL}}$ projections is reversed, which presents an additional complication for establishing convergence to the mirror Schrödinger bridge. To overcome this challenge, we make use of an observation made by Vargas & Nüsken (2023, section 4.1 and proof of Proposition 4.1): in traditional IPFP, we can reverse one or both of the direct $D_{\mathrm{KL}}$ projections (2) while preserving the sequence of iterates obtained. In particular, they prove:
Lemma 2.
Let $\pi_{0},\,\pi_{1}$ be probability distributions on $\mathbb{R}^{n}$, and let ${\mathbb{P}}\in\mathcal{P}(C([0,1],\mathbb{R}^{n}))$ be any path measure. Then we have the following identities relating direct to reverse $D_{\mathrm{KL}}$ projections:
$\operatorname*{arg\,min}_{{\mathbb{Q}}\in{\mathbb{D}}(\cdot,\pi_{1})}D_{\mathrm{KL}}\left({\mathbb{Q}}\;\|\;{\mathbb{P}}\right)=\operatorname*{arg\,min}_{{\mathbb{Q}}\in{\mathbb{D}}(\cdot,\pi_{1})}D_{\mathrm{KL}}\left({\mathbb{P}}\;\|\;{\mathbb{Q}}\right)$
$\operatorname*{arg\,min}_{{\mathbb{Q}}\in{\mathbb{D}}(\pi_{0},\cdot)}D_{\mathrm{KL}}\left({\mathbb{Q}}\;\|\;{\mathbb{P}}\right)=\operatorname*{arg\,min}_{{\mathbb{Q}}\in{\mathbb{D}}(\pi_{0},\cdot)}D_{\mathrm{KL}}\left({\mathbb{P}}\;\|\;{\mathbb{Q}}\right).$
Using Lemma 2, we obtain the following result, which states that the Schrödinger bridge can be equivalently defined in terms of reverse $D_{\mathrm{KL}}$. We defer the proof to Appendix A.
Proposition 3.
Let $\pi_{0},\,\pi_{1}$ be probability distributions on $\mathbb{R}^{n}$ and let ${\mathbb{Q}}^{0}\in\mathcal{P}(C([0,1],\mathbb{R}^{n}))$ be any path measure. Then the Schrödinger bridge ${\mathbb{Q}}_{\rm{SB}}$ with respect to ${\mathbb{Q}}^{0}$ is the unique solution to the following pair of optimization problems:
@@ 4.2 Convergence
${\mathbb{Q}}_{\rm{SB}}=\operatorname*{arg\,min}_{{\mathbb{Q}}\in{\mathbb{D}}(\pi_{0},\pi_{1})}D_{\mathrm{KL}}\left({\mathbb{Q}}\;\|\;{\mathbb{Q}}^{0}\right)=\operatorname*{arg\,min}_{{\mathbb{Q}}\in{\mathbb{D}}(\pi_{0},\pi_{1})}D_{\mathrm{KL}}\left({\mathbb{Q}}^{0}\;\|\;{\mathbb{Q}}\right).$
We are now ready to use Lemma 2 and Proposition 3 to prove Theorem 1.
Proof of Theorem 1.
Equipped with Lemma 2, we can reverse the $D_{\mathrm{KL}}$ projections in the steps given by (4). Then we obtain the following sequence of pairs of reverse $D_{\mathrm{KL}}$ projections:
${\mathbb{P}}^{2k+1}=\operatorname*{arg\,min}_{{\mathbb{P}}\in{\mathbb{D}}(\cdot,\pi)}D_{\mathrm{KL}}\left({\mathbb{P}}^{2k}\;\|\;{\mathbb{P}}\right)$ (reverse $D_{\mathrm{KL}}$ projection)
${\mathbb{P}}^{2k+2}=\operatorname*{arg\,min}_{{\mathbb{P}}\in{\mathbb{S}}}D_{\mathrm{KL}}\big{(}{{\mathbb{P}}^{2k+1}}\;\|\;{\mathbb{P}}\big{)},$ (reverse $D_{\mathrm{KL}}$ projection)
We apply the Pythagorean theorem for reverse $D_{\mathrm{KL}}$ projections, which shows that
@@
$D_{\mathrm{KL}}\left({\mathbb{P}}^{0}\;\|\;{\mathbb{P}}_{\rm{MSB}}\right)=\sum_{i=1}^{\infty}D_{\mathrm{KL}}\left({\mathbb{P}}^{i-1}\;\|\;{\mathbb{P}}^{i}\right)+\lim_{k\to\infty}D_{\mathrm{KL}}\left({\mathbb{P}}^{k}\;\|\;{\mathbb{P}}_{\rm{MSB}}\right)$ (6)
Since KL divergences are always nonnegative, the sequence of partial sums in (6) is nondecreasing and bounded, so the sum must converge. Thus, for any $\epsilon>0$, we can choose $N$ sufficiently large to ensure that $D_{\mathrm{KL}}\left({\mathbb{P}}^{n_{1}}\;\|\;{\mathbb{P}}^{n_{2}}\right)\leq\epsilon$ for all $n_{2}>n_{1}>N$. By Pinsker’s Inequality, we have that the same property holds with $D_{\mathrm{KL}}$ replaced by $\delta_{\rm{TV}}$, i.e., the metric induced by total variation. Thus, the sequence ${\mathbb{P}}^{k}$ is Cauchy with respect to $\delta_{\rm{TV}}$. Since the space of path measures is complete with respect to this metric, there exists a limit ${\mathbb{P}}^{k}\to{\mathbb{P}}^{\star}$. But just as we argued in the proof of Proposition 3, we can show that $D_{\mathrm{KL}}\left({\mathbb{P}}^{0}\;\|\;{\mathbb{P}}_{\rm{MSB}}\right)=D_{\mathrm{KL}}\left({\mathbb{P}}^{0}\;\|\;{\mathbb{P}}^{\star}\right)$, so by uniqueness of the Schrödinger bridge with respect to reverse $D_{\mathrm{KL}}$, as shown in Proposition 3, it follows that ${\mathbb{P}}_{\rm{MSB}}={\mathbb{P}}^{\star}$. ∎

### v2

@@ 4.1 Symmetrized Iterative Proportional Fitting
Take the reference path measure ${\mathbb{P}}^{0}$ to be time-symmetric. As an example, we can take ${\mathbb{P}}^{0}$ to be associated to an Ornstein–Uhlenbeck process ${\mathbf{X}}_{t}$ (at equilibrium) given by an SDE of the form ${\rm{d}}{\mathbf{X}}_{t}=-\alpha{\mathbf{X}}_{t}{\rm{d}}t+\sigma{\rm{d}}{\mathbf{W}}_{t}$, for $\alpha>0$, or more generally any reversible diffusion process. We consider the following iterative scheme involving a pair of direct $D_{\mathrm{KL}}$ projection steps:
${\mathbb{P}}^{2k+1}=\operatorname*{arg\,min}_{{\mathbb{P}}\in{\mathbb{D}}(\pi,\cdot)}D_{\mathrm{KL}}\left({\mathbb{P}}\;\|\;{\mathbb{P}}^{2k}\right)$ (direct) (4)
${\mathbb{P}}^{2k+2}=\operatorname*{arg\,min}_{{\mathbb{P}}\in{\mathbb{S}}}D_{\mathrm{KL}}\big{(}{\mathbb{P}}\;\|\;{{\mathbb{P}}^{2k+1}}\big{)},$ (direct) (5)
@@
where ${\mathbb{S}}$ is the set of time-symmetric path measures with no marginal constraints. This scheme is a dynamical analogue for path measures of the symmetrized version of Sinkhorn’s algorithm described by Kurras (2015). Each iteration of this scheme in (4)-(5) is designed to obtain the time-symmetric measure ${\mathbb{P}}$ that minimizes the objective while remaining close in $D_{\mathrm{KL}}$ divergence to the measure obtained in the previous half iteration, which satisfies the initial marginal constraint $\pi$. As direct $D_{\mathrm{KL}}$ projections satisfy the Pythagorean theorem for $D_{\mathrm{KL}}$ divergences, we have the following result:
Lemma 4.1.
The scheme in (4)-(5) converges to the mirror Schrödinger bridge.
This is proven by Ruschendorf (1995). A theoretical requirement for this scheme is that the reference measure $\mathbf{P}^{0}$ be time-symmetric, so standard Brownian motion cannot be used as a prior. As we show in §4.2, the first projection step (4) has a simple analytical solution, but the problem of computing the direct $D_{\mathrm{KL}}$ projection (5) onto the set ${\mathbb{S}}$ of time-symmetric path measures is considerably more difficult, and it remains open as to whether this projection admits an explicit analytical description, like for the projection (4). Thus, to obtain an algorithmic improvement on IPFP that leverages the symmetry of our problem, we would like a more efficient alternative to explicitly solving (5).
4.2 Alternating Minimization Procedure (AMP)
A key insight of our paper is that the symmetrization step simplifies dramatically if we reverse the direction of the $D_{\mathrm{KL}}$ divergence. This motivates us to consider a modified version of our symmetrized IPFP scheme, where the $D_{\mathrm{KL}}$ divergence in (5) is reversed. We prove in this section that the resulting alternating minimization scheme (1) has a simple analytical solution for the symmetrization step, and (2) agrees with symmetrized IPFP to first order.
More precisely, take the reference path measure ${\mathbb{P}}^{0}$ to be time-symmetric. As an example, we can take ${\mathbb{P}}^{0}$ to be associated to an Ornstein–Uhlenbeck process ${\mathbf{X}}_{t}$ given by an SDE of the form ${\rm{d}}{\mathbf{X}}_{t}=-\alpha{\mathbf{X}}_{t}{\rm{d}}t+\sigma{\rm{d}}{\mathbf{W}}_{t}$, for $\alpha>0$, or more generally any reversible diffusion process. We propose the following iterative scheme:
${\mathbb{P}}^{2k+1}=\operatorname*{arg\,min}_{{\mathbb{P}}\in{\mathbb{D}}(\pi,\cdot)}D_{\mathrm{KL}}\left({\mathbb{P}}\;\|\;{\mathbb{P}}^{2k}\right)$ (direct) (6)
${\mathbb{P}}^{2k+2}=\operatorname*{arg\,min}_{{\mathbb{P}}\in{\mathbb{S}}}D_{\mathrm{KL}}\big{(}{{\mathbb{P}}^{2k+1}}\;\|\;{\mathbb{P}}\big{)},$ (reverse) (7)
where ${\mathbb{S}}$ is the set of time-symmetric path measures with no marginal constraints. This scheme is an instance of AMP and differs from IPFP in that it alternates between direct and reverse $D_{\mathrm{KL}}$ projections. To see this, note that (6) is a direct $D_{\mathrm{KL}}$ projection and coincides with the odd-numbered steps in the IPFP iterations (2), whereas (7) is a reverse $D_{\mathrm{KL}}$ projection, as the KL divergence is being computed against the optimization parameter ${\mathbb{P}}$ instead of the previously produced path measure ${\mathbb{P}}^{2k+1}$.
If for some reason ${\mathbb{P}}^{2k+1}$ is not supported at a point, then the solution to the minimization in (7) may not be unique, in which case we define ${\mathbb{P}}^{2k+2}$ to be any minimizer. But this is no problem: in what follows, we will prove Propositions 4.2 and 4.3, which together imply that the subsequent odd-numbered iterate ${\mathbb{P}}^{2k+3}$ is always uniquely determined, regardless of the choice of ${\mathbb{P}}^{2k+2}$. We now explore each of the projections in our alternating minimization scheme in detail.
Projection (6). We can compute the direct $D_{\mathrm{KL}}$ projection onto the set of path measures with a prescribed initial marginal distribution $\pi$ following the trajectory-caching method developed and applied by Vargas et al. (2021); De Bortoli et al. (2021). Let $\pi$ be a probability distribution on $\mathbb{R}^{n}$, and let ${\mathbb{P}}\in{\mathbb{D}}(\pi,\cdot)$ and ${\mathbb{P}}^{\dagger}\in{\mathbb{S}}$ be path measures corresponding to diffusion processes. Write $f_{t}^{{\mathbb{P}}}$ and $b_{t}^{{\mathbb{P}}}$ for the forward and backward drift functions corresponding to ${\mathbb{P}}$, and write $v_{t}^{{\mathbb{P}}^{\dagger}}=f_{t}^{{\mathbb{P}}^{\dagger}}=b_{t}^{{\mathbb{P}}^{\dagger}}$ for the drift of ${\mathbb{P}}^{\dagger}$. As a consequence of Girsanov’s theorem, we can write $D_{\mathrm{KL}}({\mathbb{P}}\;\|\;{\mathbb{P}}^{\dagger})$ explicitly in terms of $f_{t}^{{\mathbb{P}}}$ and $v_{t}^{{\mathbb{P}}^{\dagger}}$, or equivalently in terms of $b_{t}^{{\mathbb{P}}}$ and $v_{t}^{{\mathbb{P}}^{\dagger}}$; for references, see (Chen et al., 2016, §3) as well as (Winkler et al., 2023, §2.2, §2.3). Indeed, for some constants $C_{1},\,C_{2}$, we have
@@ 4.2 Alternating Minimization Procedure (AMP)
$D_{\mathrm{KL}}({\mathbb{P}}\;\|\;{\mathbb{P}}^{\dagger})$ $=C_{1}+\frac{1}{2\sigma^{2}}\int_{0}^{1}\mathbb{E}_{{\mathbb{P}}}\left[\left(f_{t}^{{\mathbb{P}}}({\mathbf{X}}_{t})-v_{t}^{{\mathbb{P}}^{\dagger}}({\mathbf{X}}_{t})\right)^{2}\right]{\rm{d}}t$ (8)
$=C_{2}+\frac{1}{2\sigma^{2}}\int_{0}^{1}\mathbb{E}_{{\mathbb{P}}}\left[\left(b_{t}^{{\mathbb{P}}}({\mathbf{X}}_{t})-v_{t}^{{\mathbb{P}}^{\dagger}}({\mathbf{X}}_{t})\right)^{2}\right]{\rm{d}}t.$ (9)
In light of (8) and (9), and because drift functions are more amenable to modeling and estimation than path measures, it is convenient to recast the steps of our AMP scheme as iterative computations of drifts associated to $D_{\mathrm{KL}}$ projections. The following result is an immediate consequence of (8):
Proposition 4.2.
The direct $D_{\mathrm{KL}}$ projection of ${\mathbb{P}}^{\dagger}$ onto the space ${\mathbb{D}}(\pi,\cdot)$ is given by the unique path measure ${\mathbb{P}}$ with initial marginal $\pi$ and forward drift $f_{t}^{{\mathbb{P}}}$ equal to the drift $v_{t}^{{\mathbb{P}}^{\dagger}}$ of ${\mathbb{P}}^{\dagger}$.
In our AMP scheme, we employ Proposition 4.2 by taking ${\mathbb{P}}={\mathbb{P}}^{2k+1}\in{\mathbb{D}}(\pi,\cdot)$ to have drift equal to that of ${\mathbb{P}}^{\dagger}={\mathbb{P}}^{2k}$ for each $k\geq 0$.
Projection (7). As before, we are interested in computing the associated time-symmetric drift, rather than the path measure itself. To this end, let $\pi$ be a probability distribution on $\mathbb{R}^{n}$, and let ${\mathbb{P}}\in{\mathbb{D}}(\pi,\cdot)$ and ${\mathbb{P}}^{\dagger}\in{\mathbb{S}}$ be path measures corresponding to diffusion processes. Suppose we seek to minimize $D_{\mathrm{KL}}({\mathbb{P}}\;\|\;{\mathbb{P}}^{\dagger})$ over all ${\mathbb{P}}^{\dagger}\in{\mathbb{S}}$. Using (8) and (9), we can write $D_{\mathrm{KL}}({\mathbb{P}}\;\|\;{\mathbb{P}}^{\dagger})$ explicitly in terms of the forward and backward drift functions of the SDE corresponding to the path measures ${\mathbb{P}}$ and ${\mathbb{P}}^{\dagger}$. First, we can combine (8) and (9) to rewrite $D_{\mathrm{KL}}({\mathbb{P}}\;\|\;{\mathbb{P}}^{\dagger})$ in a time-symmetric formulation as follows:
@@
$D_{\mathrm{KL}}({\mathbb{P}}\;\|\;{\mathbb{P}}^{\dagger})=C+\frac{1}{4\sigma^{2}}\int_{0}^{1}{\mathbb{E}_{{\mathbb{P}}}}\bigg{[}\big{(}f_{t}^{{\mathbb{P}}}({\mathbf{X}}_{t})-v_{t}^{{\mathbb{P}}^{\dagger}}({\mathbf{X}}_{t})\big{)}^{2}+\left(b_{t}^{{\mathbb{P}}}({\mathbf{X}}_{t})-v_{t}^{{\mathbb{P}}^{\dagger}}({\mathbf{X}}_{t})\right)^{2}\bigg{]}{\rm{d}}t,$ (10)
where $C$ is a constant. A key benefit of considering the reverse $D_{\mathrm{KL}}$ projection is that the expectations in (8) and (9) are taken with respect to the fixed path measure ${\mathbb{P}}$, and not with respect to the varying ${\mathbb{P}}^{\dagger}$. Using calculus of variations, we can then compute a closed-form expression for the drift of the minimizer of $D_{\mathrm{KL}}({\mathbb{P}}\;\|\;{\mathbb{P}}^{\dagger})$ over ${\mathbb{P}}^{\dagger}\in{\mathbb{S}}$. Note that the sum of squares inside the expectation on the right-hand side above is always nonnegative. Consequently, to minimize $D_{\mathrm{KL}}({\mathbb{P}}\;\|\;{\mathbb{P}}^{\dagger})$, it suffices to choose ${v_{t}^{{\mathbb{P}}^{\dagger}}}$ so that it minimizes this sum of squares at each time $t$. Taking the first variation of this sum with respect to $v_{t}^{{\mathbb{P}}^{\dagger}}$, setting the result equal to zero, and solving for the optimal $v_{t}^{{\mathbb{P}}^{\dagger}}$, we obtain:
In other words, we have proven the following characterization of the projection (7):
Proposition 4.3.
The reverse $D_{\mathrm{KL}}$ projection of ${\mathbb{P}}$ onto the space ${\mathbb{S}}$ is a time-symmetric path measure with drift given by the average of the forward and backward drifts of ${\mathbb{P}}$.
It is natural to ask what relation the explicit formula (11) for the minimizer of the reverse $D_{\mathrm{KL}}$ projection (7) has to the minimizer of the direct $D_{\mathrm{KL}}$ projection (5) in symmetrized IPFP. The following proposition establishes that these two minimizers agree to first order (see proof in Appendix A).
Proposition 4.4.
Let $v_{t}^{\star}$ be the drift corresponding to the minimizer of the $D_{\mathrm{KL}}$ projection (5). Then
$v_{t}^{\star}(x)=\frac{1}{2}\left(f_{t}^{{\mathbb{P}}}(x)+b_{t}^{{\mathbb{P}}}(x)\right)+O\big{(}\|f_{t}^{{\mathbb{P}}}(x)-b_{t}^{{\mathbb{P}}}(x)\|^{2}\big{)}$
For alternating minimization schemes, like (6)-(7), some convergence results were established by Csiszár and Tusnády (1984). For instance, their Theorem 3 implies that the sequence whose $(2k+1)$-th term is $D_{\mathrm{KL}}({\mathbb{P}}^{2k+1}\;\|\;{\mathbb{P}}^{2k})$ and whose $(2k+2)$-th term is $D_{\mathrm{KL}}({\mathbb{P}}^{2k+1}\;\|\;{\mathbb{P}}^{2k+2})$ monotonically decreases to a limiting value equal to
$D_{\mathrm{KL}}({\mathbb{D}}(\pi,\cdot),{\mathbb{S}})\coloneqq\inf_{\begin{subarray}{c}{\mathbb{P}}\in{\mathbb{D}}(\pi,\cdot)\\ {\mathbb{P}}^{\dagger}\in{\mathbb{S}}\end{subarray}}D_{\mathrm{KL}}({\mathbb{P}}\;\|\;{\mathbb{P}}^{\dagger}).$
But this limiting value is zero, as the intersection ${\mathbb{D}}(\pi,\cdot)\cap{\mathbb{S}}$ is nonempty. Thus, the sequence of iterates ${\mathbb{P}}^{k}$ obtained by our AMP actually converges to a time-symmetric path measure with desired marginals (in situations where alternating projections are performed onto nonintersecting sets, convergence is much trickier to establish; e.g., for discrete probability measures, it was proven by Csiszár and Tusnády (1984) that the odd-numbered iterates converge, but the even-numbered ones may not).
The limit of our alternating scheme, while it exists, may not necessarily coincide with that of IPFP, because direct $D_{\mathrm{KL}}$ projections encourage mean-seeking behavior, whereas reverse $D_{\mathrm{KL}}$ projections encourage mode-seeking behavior. This concurs with an observation of Kurras (2015) in the static case, where they note that taking the arithmetic mean of the transport matrix and its transpose to perform the symmetrization step can cause the method to converge to a limit different from the optimal one. So far, we proved that our reversed symmetrization update agrees with the direct $D_{\mathrm{KL}}$ projection to first order. Thus, we expect that our alternating scheme converges to a good approximation of the Schrödinger bridge. We confirm this expectation in two different ways. First, we show that the limit of our approximate scheme is close to the Schrödinger bridge in the Gaussian case and is numerically indistinguishable from the result of IPFP; see Fig. 2 and 6. Second, we formulate reasonable hypotheses under which the difference between the mirror Schrödinger bridge and the limit of our alternating scheme, i.e., the mirror bridge, can be explicitly bounded; see Appendix B.